Let P be a point on the parabola y-2y-4x+5=0, such that the tangent on the parabola at P intersects the directrix at point Q. let R be the point that divides the line segment QP externally in the ratio ½ : 1 . Find the locus of R.

Grade:10

1 Answers

Vikas TU
6 years ago
Dear Student,
vertex of parabola = 1,1
       Directrix = x=0
       Eq of parabola : (y-1)^2 = 4(x-1) ; a=1
        P= (t^2,2t)
        Q=(0,t)
        R= ( 2t^2, 3t)
        For point R, x=2t^2 ; y= 3t
        Eliminating t from the two equations, we get 9x= 2y^2 as the locus of R 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)